### v2

# MrGuard: A Multilingual Reasoning Guardrail for Universal LLM Safety

Abstract
Large Language Models (LLMs) are susceptible to adversarial attacks such as jailbreaking, which can elicit harmful or unsafe behaviors. This vulnerability is exacerbated in multilingual settings, where multilingual safety-aligned data is often limited. Thus, developing a guardrail capable of detecting and filtering unsafe content across diverse languages is critical for deploying LLMs in real-world applications. In this work, we introduce a multilingual guardrail with reasoning for prompt classification. Our method consists of: (1) synthetic multilingual data generation incorporating culturally and linguistically nuanced variants, (2) supervised fine-tuning, and (3) a curriculum-based Group Relative Policy Optimization (GRPO) framework that further improves performance. Experimental results demonstrate that our multilingual guardrail, MrGuard, consistently outperforms recent baselines across both in-domain and out-of-domain languages by more than 15%. We also evaluate MrGuard’s robustness to multilingual variations, such as code-switching and low-resource language distractors in the prompt, and demonstrate that it preserves safety judgments under these challenging conditions. The multilingual reasoning capability of our guardrail enables it to generate explanations, which are particularly useful for understanding language-specific risks and ambiguities in multilingual content moderation.
Warning: This paper contains potentially harmful examples.
MrGuard: A Multilingual Reasoning Guardrail for Universal LLM Safety
Yahan Yang University of Pennsylvania yangy96@seas.upenn.edu                        Soham Dan Microsoft sohamdan@microsoft.com                        Shuo Li University of Pennsylvania lishuo1@seas.upenn.edu
Dan Roth University of Pennsylvania danr@seas.upenn.edu                        Insup Lee University of Pennsylvania lee@seas.upenn.edu
1 Introduction
Figure 1: The guardrails specialized in English (GuardR, (Liu et al., 2025)) are providing different predictions for English and Chinese inputs with the same semantic meaning. Our MrGuard can analyze the Chinese prompts with explanation and provide correct safety prediction.
Large Language Models (LLMs) have demonstrated impressive capabilities in cross-lingual knowledge transfer, enabling them to perform a variety of tasks across multiple languages even when fine-tuned on primarily monolingual datasets (Touvron et al., 2023; Brown et al., 2020; Qin et al., 2024). This cross-lingual ability is largely attributed to their large-scale and diverse pretraining corpora, which allow LLMs to handle multilingual inputs without requiring significant multilingual data for downstream task adaptation. LLMs are increasingly being applied in a wide range of real-world applications, including conversational agents, educational tools, and medical assistants. However, despite these advancements, current LLMs are not yet robust or reliable enough for deployment in safety-critical environments. They can be intentionally misused to promote harmful behavior, generate offensive or biased content, or even bypass safety mechanisms through adversarial prompting (i.e., jailbreaking) (Andriushchenko et al., 2024; Chao et al., 2023). These vulnerabilities are further amplified in multilingual settings, particularly for low-resource languages, where models may lack proper safety alignment due to limited training signals or evaluation benchmarks (Deng et al., 2023; Wang et al., 2023).
To address these issues, safety alignment techniques—such as Reinforcement Learning from Human Feedback (RLHF) and Direct Preference Optimization (DPO)—have been proposed to align LLM behavior with human values and reduce unsafe generations (Ouyang et al., 2022; Rafailov et al., 2023). Another line of work focuses on building standalone safety classifiers or guardrails, which act as filters to detect and block unsafe user prompts or model generations without modifying the LLM itself (Inan et al., 2023; Ghosh et al., 2024). These lightweight safety modules are advantageous in being more efficient and easier to deploy or update (Table 1). Most existing methods are primarily English-centric (Liu et al., 2025; Kang and Li, 2024b; Yuan et al., 2024) which cannot handle multilingual content moderation (Yang et al., 2024). As shown in Figure 1, the guardrail model successfully identifies the unsafe user prompt in English but fails to detect its semantically equivalent counterpart in Chinese. Moreover, without explanations, it becomes difficult to understand the rationale behind the guardrail’s decisions [^1] We interchangeably use guard and guardrail through the paper..
Base Model Data R
GuardR
(Liu et al., 2025)
LlaMa-3.1-8B 127k EN Yes
DUO-Guard
(Deng et al., 2025)
QWEN-0.5B
1679k EN
100k MUL
No
Aegis-2.0
(Ghosh et al., 2025)
LlaMa-3.1-8B
-Instruct
30k EN No
LlaMa-Guard-3
(Inan et al., 2023)
LlaMa-3.1-8B Unknown No
WildGuard
(Han et al., 2024)
Mistral-7B 86.8K EN No
MrGuard (Ours)
LlaMa-3.1-8B
-Instruct
30k EN
6k MUL
Yes
Table 1: Configurations of recent state-of-the-art guardrails. Base Model refers to the underlying language model used by each guardrail. Data indicates the dataset used for training the guardrail, where EN denotes English-only data and MUL refers to multilingual (non-English) data. R specifies whether the guardrail is trained with reasoning capability.
To bridge this gap, our work is the first one to focus on building a guardrail tailored for multilingual safety scenarios with reasoning ability. We aim to design a robust, reasoning-aware safety guardrail that can effectively moderate harmful prompts across diverse languages and cultural contexts. Our contributions can be listed as follows:
  • We introduce MrGuard, a multilingual reasoning-enhanced guardrail for prompt moderation that improves performance and robustness across languages. Our approach combines curriculum learning (Bengio et al., 2009a) with Group Relative Policy Optimization (GRPO) (Shao et al., 2024b) to gradually introduce more culturally diverse variants at the post-training stage.
  • We achieve state-of-the-art results on several multilingual safety benchmarks, outperforming all baselines in prompt classification accuracy. We further demonstrate MR.Guard’s robustness on multilingual variations such as code-switching and sandwich Upadhayay and Behzadan (2024) attacks. Our results show that post-training with reasoning abilities significantly improves the robustness and performance of guardrails on multilingual prompt classification.
  • We present a comprehensive evaluation of our reasoning-enhanced guardrail with key metrics such as cross-lingual consistency and reasoning fidelity which establishes a strong baseline for future assessments of guardrail reasoning capabilities.
2 Related Work
2.1 Multilingual LLM Safety
While LLMs demonstrate strong cross-lingual capabilities on multilingual downstream tasks, their ability to handle unsafe content in multilingual settings remains largely unknown, and there is still significant room for improving their robustness to multilingual inputs. Prior studies (Wang et al., 2023; Deng et al., 2023) have shown that LLMs are vulnerable to non-English jailbreaking prompts, especially in low-resource languages. Follow-up work (Yoo et al., 2024) uses GPT-4 to combine parallel jailbreaking queries in Deng et al. (2023) from different languages into a single code-switching prompt, demonstrating that such prompts further increase the attack success rate compared to monolingual attacks. Recent work (de Wynter et al., 2024; Jain et al., 2024; Ye et al., 2023) collects multilingual moderation datasets to investigate the ability of LLMs to respond to multilingual harmful prompts/responses and assess whether guardrails effectively filter them out. They all show that existing guard or encoder-only classifiers cannot adequately handle multilingual content moderation. Upadhayay and Behzadan (2024) has introduced an attack against LLMs by embedding jailbreaking prompts within unrelated, low-resource, but safe inputs. Their results show that both proprietary and open-source models are vulnerable to these low-resource distractors, often following the embedded unsafe instructions and generating harmful content. These findings all underscore the urgent need for robust guardrails capable of detecting and mitigating unsafe behavior in multilingual settings.
2.2 Guardrails for safeguarding LLMs
Recent guardrail research (Li et al., 2024; Inan et al., 2023; Rebedea et al., 2023; Ghosh et al., 2025; Kang and Li, 2024a) have leveraged pre-trained small language models (SLMs), such as LlaMa-2/3.1-7B (Touvron et al., 2023) and Mistral-7B (Jiang et al., 2023), to distinguish between safe and unsafe content. These methods have demonstrated promising results in detecting harmful inputs in English compared to encoder-only models. Yuan et al. (2024) has enhanced base guardrail models by incorporating energy-based data generation and combining guardrail predictions with k-nearest neighbors (kNN) predictions. Additionally, Kang and Li (2024b) has introduced a knowledge-based logical reasoning framework, which first asks the model to determine whether an input belongs to a predefined risk category and then uses a probabilistic graphical model to estimate the likelihood of unsafety. GuardReasoner (Liu et al., 2025) has improved interpretability and performance by training the base model on reasoning-augmented data and applying reinforcement learning (RL) algorithm DPO (Rafailov et al., 2023) to select difficult examples. However, these efforts largely focus on English. To address multilingual safety, Deng et al. (2025) has introduced an RL-based method for generating synthetic multilingual data by iteratively and jointly updating a synthetic data generator and a guardrail model. Yet, their work targets only high-resource languages close to English. In contrast, we focus on building MrGuard: a multilingual guardrail capable of handling cultural nuances and language-specific challenges spanning languages from several different families.
To enhance the reasoning capabilities of language models, we integrate curriculum learning with a reinforcement learning (RL) algorithm known as Group Relative Policy Optimization (GRPO) Shao et al. (2024b). GRPO demonstrates superior performance compared to offline methods such as Direct Preference Optimization (DPO) (Rafailov et al., 2023), while offering improved computational efficiency over on-policy algorithms like Proximal Policy Optimization (PPO). A more comprehensive discussion of these RL approaches and their relationship to curriculum learning is provided in Appendix A.
3 Multilingual Guard with Reasoning
Figure 2: Workflow of our three-stage approach: (1) synthetic data generation, (2) supervised fine-tuning, and (3) curriculum-based Group Relative Policy Optimization (GRPO). The upper part illustrates the generation of multilingual translations and reasoning from English seed data using LLMs, followed by supervised fine-tuning. The lower part shows the construction of a curriculum by generating multilingual data with varying difficulty levels, which are then used to train the model via GRPO.
We detail our algorithm for building MrGuard: a guardrail with reasoning capabilities for multilingual content moderation in this section. The approach consists of three key components as shown in Figure 2:
1. Synthetic Data Generation – Inspired by Liu et al. (2025), we collect the analysis of safety for our seed data from a more powerful proprietary model, GPT-4o-mini [^2] We utilize GPT-4o-mini (gpt-4o-mini-2024-07-18) through all data generation stage, referred to as GPT for simplcity, in the next subsection. and use the collected data to train our model. We additionally generate multilingual data and corresponding multilingual analysis using GPT-4o-mini.
2. Supervised Fine-Tuning – We fine-tune an instruction-optimized model on the generated data to enable multilingual reasoning capabilities on content moderation, and safety classification.
3. Curriculum-Based Optimization – We combine a three-stage curriculum learning framework and GRPO Shao et al. (2024b) to align the model with desired multilingual moderation behavior.
3.1 Synthetic Data Generation
We consider an English safety training dataset, $D=\{(p_{i}^{l_{0}},y_{i})\}_{i=1}^{N}$, where $p_{i}^{l_{0}}$ is an English-language prompt and $y_{i}\in\{\text{Safe},\text{Unsafe}\}$ is its corresponding safety label. For each prompt $p_{i}^{l_{0}}$, we prompt GPT to generate reasoning for why it is labeled as $y_{i}$. This yields an augmented dataset with model-generated reasoning, denoted as
$D^{l_{0}}=\{(p_{i}^{l_{0}},e_{i}^{l_{0}},y_{i})\}_{i=1}^{N}.$
Next, we subsample a smaller set from the original dataset $D$, forming a subset $D_{\text{sub}}$. For each target language $l_{k}$, we prompt GPT to translate each English prompt $p_{i}^{l_{0}}\in D_{\text{sub}}$ into the target language, resulting in $p_{i}^{l_{k}}$. We assume that the safety label $y_{i}$ is preserved across translations. To further ensure label consistency, we prompt GPT to reassess the safety of each translated prompt $p_{i}^{l_{k}}$, If the reassessed label conflicts with the original $y_{i}$, the corresponding example is discarded from the training set. We then prompt GPT with $(p_{i}^{l_{k}},y_{i})$ to generate the corresponding reasoning $e_{i}^{l_{k}}$ in language $l_{k}$ and English $e_{i}^{l_{0}}$, yielding the dataset $D^{l_{k}}=\{(p_{i}^{l_{k}},e_{i}^{l_{0}},e_{i}^{l_{k}},y_{i})\}_{i=1}^{n}.$. Note that $D^{l_{k}}$ contains significantly fewer examples than $D^{l_{0}}$. Given $K$ target languages $\{l_{1},\ldots,l_{K}\}$, we obtain the multilingual dataset:
$D^{\text{multi}}=\{D^{l_{0}},D^{l_{1}},\ldots,D^{l_{K}}\}.$
3.2 Supervised Fine-tuning
We perform supervised fine-tuning of the base model, denoted by $\pi$, using the multilingual dataset $D^{\text{multi}}$, to enable the model to identify safe and unsafe prompts along with their reasoning. Given a data point $(p_{i}^{l_{k}},e_{i}^{l_{0}},e_{i}^{l_{k}},y_{i})$, we fine-tune the base model by applying cross-entropy loss on the tokens corresponding to both reasoning trajectories and the safety label. This enables the model to leverage the strong generalization capabilities of English while simultaneously developing multilingual reasoning skills, thereby preparing it for the subsequent reinforcement learning stage. We denote the resulting fine-tuned model as $\pi_{\text{sft}}$.
3.3 Curriculum-based GRPO
In this stage, we employ reinforcement learning to further enhance detection performance by eliciting stronger reasoning capabilities. We begin by re-sampling a subset $D^{l_{0}^{\prime}}$ from the original English safety training dataset $D^{l_{0}}$. Each prompt in $D^{l_{0}^{\prime}}$ is then translated into the target languages $l_{k}$, for $k\in\{1,\ldots,K\}$.We then introduce a curriculum-based training schedule. The intuition is that, since the base model is initially fine-tuned on an English-dominant corpus, it is more familiar with English-specific nuances, such as slang and native expressions. To guide the model in progressively learning to handle other languages as second languages, we propose a curriculum that gradually introduces more challenging native multilingual variants. These variants are derived from English sentences and are incorporated stage by stage to support step-wise multilingual adaptation. To construct the curriculum, we introduce a novel difficulty function $\operatorname{Diff}$ that quantifies the difficulty of prompts in various target languages. Specifically, all the English prompts are assigned a baseline difficulty level of 0. For a prompt $p^{l_{k}}$ in language $l_{k}$ and its corresponding English prompt $p^{l_{0}}\in D^{l_{0}^{\prime}}$, we use the prompt template shown in Figure 6 to instruct GPT to generate two challenging variants, $p^{l_{k}\prime}$ and $p^{l_{k}\prime\prime}$, enriched with slang, references to local places, institutions, foods, and other culturally or linguistically specific elements. A translation model $\pi_{\text{bt}}$ is then used to back-translate $p^{l_{k}\prime}$ and $p^{l_{k}\prime\prime}$ into English. The semantic similarity between the back-translated prompt and the original English prompt $p^{l_{0}}$ is computed using the cosine similarity function $\operatorname{cos}$. The difficulty of a back-translated prompt $p\in\{p^{l_{k}},p^{l_{k}\prime},p^{l_{k}\prime\prime}\}$ is defined as:
$\operatorname{Diff}(p)=\begin{cases}0,&\operatorname{cos}(\pi_{\text{bt}}(p),p^{l_{0}})>t_{1},\\ 1,&\operatorname{cos}(\pi_{\text{bt}}(p),p^{l_{0}})\in(t_{2},t_{1}],\\ 2,&\text{otherwise},\end{cases}$
where $t_{1}$ and $t_{2}$ are threshold hyperparameters. During training, prompts with difficulty level 0 are introduced in the first epoch. Prompts with levels 1 and 2 are progressively added in the second and third epochs, respectively, following the curriculum learning schedule.
After developing the curriculum, we apply GRPO to optimize the reference model $\pi_{\text{sft}}$ (Shao et al., 2024b). We utilize rule-based reward functions, with the following components:
Format reward ($\mathcal{R}_{f}$):
This reward penalizes formatting errors. If the output does not contain a properly formatted safety prediction (i.e., "Safety: safe" or "Safety: unsafe") which often happens in multilingual generation, the reward is $-1$. Otherwise, the reward is $1$.
Correctness reward ($\mathcal{R}_{c}$):
If the safety prediction is correct, the reward is $1$, otherwise, the reward is $-1$.
Uncertainty reward ($\mathcal{R}_{u}$):
We train an auxiliary encoder-only model $\pi_{u}$ to use the reasoning to decide whether the input is safe or not (binary classification) and take the softmax score as the reward.
$\mathcal{R}_{u}=\begin{cases}\pi_{u}(q,\hat{e}),&\text{if }\text{prediction iscorrect}\\ -\pi_{u}(q,\hat{e})&\text{if }\text{prediction is incorrect}\end{cases}$
Language reward ($\mathcal{R}_{lang}$):
For the second and third stages, the input sentences are more native to the target language. We hypothesize that language-specific reasoning enhances the model’s understanding in this setting. To encourage the model to generate reasoning in the target language, we add this language reward,
$\mathcal{R}_{lang}=\begin{cases}0.5,&\text{if }\text{difficulty}=1\\ 1.0,&\text{if }\text{difficulty}=2\\ 0.0,&\text{otherwise}\end{cases}$
Finally, the individual reward signals are combined linearly to produce a single scalar reward value:
$\mathcal{R}=\mathcal{R}_{f}+\mathcal{R}_{c}+\mathcal{R}_{u}+\mathcal{R}_{lang}.$
With the reward signals defined, we apply the original GRPO algorithm to optimize the reference model $\pi_{\text{sft}}$. For a detailed explanation of GRPO, please refer to Appendix E and Shao et al. (2024a).
4 Experiments
4.1 Experimental Setup
In our experiments, we use the training set from Aegis-2.0-Safety (Ghosh et al., 2025) as the English seed data. Our base model is LLaMA-3.1-8B-Instruct (Aaron Grattafiori, 2024), and we apply QLoRA (Dettmers et al., 2023) for parameter-efficient fine-tuning during both the SFT and GRPO stage [^3] More details on the training hyperparameters and configurations are in Appendix B.. To construct the curriculum using back-translation, we employ the facebook/nllb-200-3.3B model for translation and use all-MiniLM-L6-v2 to compute the sentence embeddings of both the original and the back-translated sentences. For the difficulty threshold, we set $t_{1}=0.85$ and $t_{2}=0.7$. To determine the language of the sampled output, we utilize an xlm-based language detector  [^4] The full name of the language detector: papluca/xlm-roberta-base-language-detection. Our experiments divide the test sets into two categories: in-domain languages, which are covered during training, and out-of-domain languages, which are not seen during training.
Models RTP-LX Aya XSafety Wildchat MultiJail
ID OOD ID OOD ID OOD ID OOD ID OOD
DUO-Guard 61.02 45.61 58.58 44.50 66.62 61.68 56.89 67.29 73.60 30.20
GuardR 78.47 61.35 90.14 90.30 83.12 80.80 80.01 82.56 91.89 78.29
LlaMa-Guard-3 45.78 45.91 79.15 82.01 61.87 60.89 67.36 69.50 78.91 75.25
Aegis-2.0 53.52 37.16 43.96 38.93 40.85 26.10 60.35 62.51 52.98 19.08
Wildguard 65.39 34.28 74.57 77.01 74.36 60.63 67.91 64.76 74.81 50.77
MrGuard (Ours) 90.95 86.90 98.18 98.11 94.56 92.06 91.45 92.68 97.41 96.15
Table 2: Performance of different guardrails to identify multilingual safety across five benchmark datasets. We report F1 scores as the evaluation metric and bold the best-performing results for each dataset where ID refers to in-domain languages and OOD refers to out-of-domain languages. Top are baselines and the bottom part is MrGuard (ours). The model size and training dataset size are listed in Table 1.
4.2 Multilingual Content Moderation
Benchmark:
Our experiments cover 5 recent multilingual safety benchmarks: PTP_wildchat (Jain et al., 2024) [^5] We only select wildchat subset. (Wildchat), RTP_LX (de Wynter et al., 2024), aya-red-teaming (Aya)(Aakanksha et al., 2024) , MultiJail (Deng et al., 2023), and XSafety (Wang et al., 2023). We define five in-domain languages—English (EN), Arabic (AR), Spanish (ES), Chinese (ZH), and Russian (RU), which are included in the training data. To further assess generalization, we also evaluate on three out-of-domain languages (listed in Table 7) that are not included in the training set but are in the test datasets. Details of the evaluation benchmark are provided in Appendix C.
Figure 3: F1 score breakdown on the RTP_LX dataset, evaluated across 8 target languages. Here EN, AR, ES, RU, and ZH are in-domain languages, and FR, HI, SW are out-of-domain languages.
Baselines:
We compare our guardrail against several recent content moderation guardrails, both with and without reasoning capabilities. The configurations and details of the different baseline models are summarized in Table 1.
Table 2 summarizes MrGuard’s performance on several multilingual moderation benchmarks. Across both in-domain and out-of-domain languages, MrGuard consistently outperforms all baselines by a substantial margin. Additionally, we observe that guardrails with reasoning capabilities (GuardR and ours) generalize better across datasets and languages, but with our approach, MrGuard achieves state-of-the-art performance in multilingual scenarios. This showcases that MrGuard effectively captures language-specific nuances, as the test datasets are naturally generated or annotated by native speakers. Note that we restrict our experiments to compact models ($\leq$ 7 B parameters) to ensure low latency and easy deployment, in line with other guardrails (see Table 1). Even at this scale, adding reasoning at the post-training stage yields large gains in multilingual classification.
Moreover, we present a language-wise breakdown in Figure 3 with more results in Appendix B. Our model consistently outperforms the baselines across a wide range of languages, although some baselines achieve comparable results on the English subset. We also observe that training on English-only data negatively impacts generalization to non-English languages. For example, although Aegis-2.0 and WildGuard perform well on English inputs, their performance degrades significantly on non-English data, whereas MrGuard maintains high performance and shows minimal performance drop, even on unseen languages. This further demonstrates the robustness and cross-lingual generalization capabilities of our model.
Furthermore, Figure 4 presents example reasoning generated by our multilingual guardrail in various languages, showing that the reasoning accurately analyzes and justifies the safety prediction of the input prompts. The reasoning can help users understand the rationale of MrGuard behind its safety decisions.
Figure 4: Example of reasoning generated from our multilingual guardrail.
4.3 Robustness to Multilingual Perturbations
In this section, we investigate the potential of using guardrails to identify unsafe prompts that involve perturbations specific to the multilingual setting. We consider two existing multilingual attacks: 1) Yoo et al. (2024) generates code-switching prompts using two parallel datasets, MultiJail and XSafety, and GPT (CSRT) 2) Sandwich attack (Upadhayay and Behzadan, 2024) (Sandwich), where jailbreaking prompts are embedded within benign prompts in lower-resource languages. Both attack strategies have demonstrated that LLMs are more vulnerable to these challenging input variants and are more likely to produce harmful responses. We crafted the variations of MultiJail and XSafety datasets using those two adversarial attacks, and examples are shown in Figure 9666The configuration and details of the generated attack are in Appendix D..
We benchmark several guardrail methods against the adversarial multilingual attacks. Table 3 and 4 reports the F1 scores before and after the attacks, along with the corresponding performance changes. As shown in the tables, all methods experience a decline in F1 score after the attack, demonstrating the effectiveness of both adversarial strategies. Notably, our method not only outperforms the baselines but also exhibits a smaller reduction in F1 score. Our experiments show that incorporating reasoning alongside safety classification significantly enhances the guardrail’s robustness against multilingual adversarial prompts.
Models EN $\uparrow$ Avg-CSRT $\uparrow$ $\Delta$$\downarrow$
DUO-Guard 90.62 71.22 19.40
GuardR 95.35 92.95 2.40
LlaMa-Guard-3 80.68 77.12 3.56
Aegis-2.0 86.69 45.59 41.10
Wildguard 95.17 81.83 13.34
MrGuard (Ours) 98.22 96.68 1.54
Table 3: F1 scores on code-switching prompts evaluated on the MultiJail datasets. The best-performing results across models are highlighted in bold. $\Delta$ represents the difference between the F1 score on English prompts and the averaged F1 score over all code-switching variants across both ID and OOD languages.
Models
Avg-
Orig
$\uparrow$
Avg-
Sandwich
$\uparrow$
$\Delta$$\downarrow$
DUO-Guard 51.90 0.58 51.32
GuardR 85.09 78.78 6.31
LlaMa-Guard-3 77.08 8.65 68.43
Aegis-2.0 36.03 2.42 33.61
Wildguard 62.79 45.57 17.22
MrGuard (Ours) 96.50 92.37 4.13
Table 4: F1 scores on sandwich attacks evaluated on the MultiJail dataset. The best-performing results across models are highlighted in bold. Avg-Orig indicates the average F1 score on before attack, and the average F1 score after sandwich attack across both ID and OOD languages. $\Delta$ represents the difference between them.
5 Discussion
In this section, we conduct a deeper analysis of our framework and results, including ablation experiments of the proposed approach, evaluation of the fidelity of reasoning and safety predictions, and cross-language consistency.
5.1 Ablation Study
In this section, we conduct an ablation study to investigate the effectiveness of GRPO and curriculum learning, and various components of the reward function to show that all of them help improve the generalization of our guardrail’s performance on different languages across different datasets.
RTP-LX Aya XSafety
Ours 89.27 98.18 93.48
wo GRPO 84.05 95.05 88.78
wo Curr 87.02 97.20 91.55
wo $\mathcal{R}_{lang},\mathcal{R}_{u}$ 88.48 97.59 92.36
wo $\mathcal{R}_{lang},\mathcal{R}_{u}$ 88.48 97.59 92.36
wo $\mathcal{R}_{lang}$ 89.55 98.35 93.07
wo $\mathcal{R}_{u}$ 88.89 97.58 92.08
Table 5: F1 scores for the ablation study. Curr denotes GRPO with curriculum learning. $\mathcal{R}_{f+a}$ represents the combination of the format reward and accuracy reward. $\mathcal{R}_{u}$ corresponds to the uncertainty reward, and $\mathcal{R}_{lang}$ denotes the language reward.
Based on the results in Table 5, we first observe that both GRPO and curriculum learning significantly improve the performance compared to $\pi_{\text{sft}}$. Consistent with prior work (DeepSeek-AI, 2025), post-training with GRPO improves the generalization across different datasets and enhances reasoning abilities. Moreover, the comparison between models trained with and without curriculum learning shows that gradually increasing the difficulty of training inputs, based on linguistic and cultural complexity, further enhances the model’s multilingual understanding. This finding underscores the value of curriculum-based learning strategies in improving robustness and generalization for multilingual safety tasks.
Furthermore, we show that all components in the reward functions positively contributes to the overall performance of MrGuard. Although removing the language reward leads to slightly better performance across different datasets, we find that the resulting model predominantly generates English reasoning. In practice, however, it is important for the guardrail to produce reasoning in the corresponding input language, making multilingual reasoning generation a valuable capability despite the marginal trade-off in accuracy. We leave the theoretical analysis of curriculum learning and reward function design of GRPO to future work.
5.2 Cross-lingual Consistency
One important characteristic for guardrails is that it assigns the same safety label to semantically equivalent prompts in different languages. To quantify this, we define the Cross-Lingual Consistency score as the fraction of parallel examples in which the model’s safety predictions agree across languages (She et al., 2024). We report consistency score on XSafety, a parallel dataset, comparing English with each target language across several models in Figure 5. From the results we observe that although our algorithm does not explicitly train for consistency, we still see improved consistency, especially for unsafe prompt classification. As shown in Figure 10, MrGuard exhibits a much smaller performance drop between ID and OOD languages compared to the other baselines.
Figure 5: Consistency score between English and non-English on the XSafety dataset, evaluated across 8 target languages. Here AR, ES, RU, and ZH are in-domain languages, and FR, HI, JA are out-of-domain languages. Higher the score the better.
5.3 Quality of Reasoning
LLMs are likely to produce hallucinations, even when guided via chain-of-thought. MrGuard is intended to help users and regulators inspect and trust its decisions, making fidelity measurement crucial. To this end, we employ a stronger LLM, GPT-4.1-mini, as a judge to automatically assess whether each explanation faithfully reflects the input and correctly drives the safety prediction. We define the Explanation Fidelity (EF) score as the fraction of reasoning sentences the judge labels as coherent out of the total sentences, and the results are shown in Table 6. Moreover, it is important to maintain the reasoning language in the same language as the prompt. We report the Language Match (LM) rate, which captures the percentage of cases where the generated reasoning is in the same language as the prompt. The results show that our reasoning is aligned with the semantics and the language of the prompt.
Lang EN AR ZH RU HI
EF 87.39 80.57 93.33 86.53 88.97
LM 97.30 98.76 99.52 98.86 99.91
Table 6: Here we report two metrics on RTP_LX dataset to evaluate the quality of the reasoning, where EF is the explanation fidelity rate and LM indicates the language matching rate. Higher score the better.
6 Conclusion
In this work, we introduce MrGuard: a multilingual reasoning-enhanced guardrail for multilingual prompt moderation. Our method consists of three key stages: synthetic data generation, supervised fine-tuning, and reinforcement learning, where we adopt GRPO with a multi-stage curriculum that progressively introduces more cultural and language-specific elements. We conduct comprehensive experiments across multiple diverse and realistic multilingual content moderation benchmarks, including challenging scenarios involving code-switching, and demonstrate that our guardrail achieves state-of-the-art performance with reasoning. We also analyze the generated reasoning to validate its reliability and ensure consistent safety preservation across languages. The reasoning ability enables multilingual users to understand the decision from MrGuard. We believe this work is an important step toward enhancing the safety of LLMs in a multilingual world.
Limitations
Language and resource coverage Due to budget and computational limits, we generated synthetic data only for high- and mid-resource languages, and relied on Aegis-2.0 as our English seed dataset. Expanding to additional seed datasets and low-resource languages could further enhance model performance and broaden the safety taxonomy to better reflect diverse user needs. Additionally, while our guardrail demonstrates strong results on both in-distribution and out-of-distribution dataset and languages, the languages represented in our evaluation remain limited.
Potential Bias We use a single LLM as a judge to verify safety labels of translations, which may introduce bias inherent to the LLM. Additionally, our current evaluation of reasoning coherence and faithfulness between explanations and final safety predictions relies on automated heuristics, which may not perfectly align with human judgments.
Ethical Statement
Our works aims to improve LLM safety for multilingual users by introducing a multilingual reasoning guardrail, which is important for building a universally reliable LLM for safety-critical applications. The generated synthetic data and models will be released, accompanied by detailed usage guidelines to prevent misuse.
Appendix A Additional Related Work
LLM Reasoning.
Several methods have been proposed to enhance the reasoning capabilities of large language models (LLMs), which can broadly be categorized into prompt engineering and post-training approaches. Prompt engineering methods, such as Chain-of-Thought (Wei et al., 2023), leverage in-context demonstrations to elicit more coherent and structured reasoning trajectories. Building on this idea, Tree-of-Thought (Yao et al., 2023) and Graph-of-Thought (Besta et al., 2024) further improve reasoning by organizing generation within tree- and graph-based logical structures. These prompt-based techniques are post-hoc in nature, enhancing reasoning without modifying the model parameters.
In contrast, post-training approaches aim to directly optimize LLMs for improved reasoning. For instance, Muennighoff et al. (2025) and Chen et al. (2023) apply supervised fine-tuning with high-quality, diverse demonstrations, while Xiong et al. (2024) utilize alignment strategies such as Direct Preference Optimization (DPO). More recently, reinforcement learning methods—including PPO and GRPO—have demonstrated strong performance in reasoning tasks (Shao et al., 2024a; DeepSeek-AI, 2025; Qwen et al., 2025; Kazemnejad et al., 2024). Among these, GRPO has gained particular attention for its superior computational efficiency compared to other reinforcement learning algorithms.
Curriculum Learning.
Training machine learning models using a progression from easy to hard examples—known as curriculum learning(Bengio et al., 2009b)—has been shown to outperform standard training approaches based on random data shuffling(Soviany et al., 2022). This paradigm has been successfully applied in both supervised learning (Graves et al., 2017; Hacohen and Weinshall, 2019; Matiisen et al., 2020) and reinforcement learning (Narvekar et al., 2020; Ren et al., 2018; Florensa et al., 2017). More recently, curriculum learning has also been explored in the context of LLM alignment (Croitoru et al., 2025).
Appendix B Experiment Setup
We used the Huggingface framework (Wolf et al., 2020) to load dataset and evaluate the guardrails and applied the default greedy decoding for all guardrails. Our training is performed on 4 NVIDIA RTX A100 (80G) GPUs, and vLLM (Kwon et al., 2023) is used to optimize inference speed.
For training data configuration, during the SFT stage, we train on the full English seed dataset (30.3K examples) combined with the translations generated from the sampled 2,000 seed examples. At the GRPO stage, we subsample another 2,000 seed examples and generate the challenging variants for curriculum learning, and we additionally include the seed English samples in the first curriculum stage to avoid losing its English ability as described in Section 3.3.
We use the TRL library (von Werra et al., 2020) for both SFT and GRPO stage. For both stages, we set the LoRA rank and alpha to 32, with a dropout rate of 0.1. During SFT, we use a learning rate of SFT is $52e-$ 5 and train for 3 epochs. For GRPO, we set the learning rate to $51e-$ 5 and the number of training epoch is 1. We conduct a hyperparameter sweep over LoRA rank and alpha values $\{8,16,32\}$, and select the best configuration based on performance on the Aegis-2.0 validation dataset. The results are the average of three training runs. All use of the packages and artifacts are consistent with their intended use and license. We used ChatGPT to refine short sentences and paragraphs and to check for grammar errors.
Appendix C Dataset Details
In our training, we translate English seed data into RU, ES, ZH, AR, so these languages are considered as our in-domain languages. We benchmark our guardrail’s performance on five multilingual datasets: PTP_wildchat (Jain et al., 2024), RTP_LX (de Wynter et al., 2024), aya-red-teaming (Aya)(Aakanksha et al., 2024) , MultiJail (Deng et al., 2023), and XSafety (Wang et al., 2023). For each dataset, we list the ID and out-of-domain (OOD) languages in Table 7. Several datasets (Aya, MultiJail, XSafety) are red-teaming datasets that consist solely of unsafe prompts intended to elicit harmful responses; we thus assume all prompts in these datasets are unsafe. For this purpose, we focus on a selected subset of topics in XSafety: Crimes and Illegal Activities, Goal Hijacking, Insult, Role Play Instruction, Unfairness and Discrimination, and Unsafe Instruction Topic. For RTP_LX, we consider prompts with an average toxicity score above 1.0 (on a scale from 1 to 5) as unsafe. For PTP_wildchat, we treat prompts with a prompt toxicity score above 0.1 (on a scale from 0 to 1) as unsafe. The language code is listed in Table 8.
RTP-LX Aya MultiJail PTP_wildchat XSafety
ID EN, AR, ES, RU, ZH-Hans EN, AR, ES, RU EN, AR, ES, ZH EN, AR, ES, RU, ZH EN, AR, ES, RU, ZH
OOD FR, HI, SW HI, FR, SR SW, IT, KO HI, FR, JA HI, FR, JA
Table 7: In-domain (ID) and out-of-domain (OOD) language coverage for each evaluation dataset.
Language Language Code
English EN
Arabic AR
Spanish ES
Russian RU
Chinese ZH
Simplified Chinese ZH-Hans
French FR
Hindi HI
Swahili SW
Serbian SR
Italian IT
Korean KO
Japanese JA
Table 8: Mapping of language to language code in the evaluation.
C.1 Generation configuration
We provide the prompts used in our experiments in this section. For synthetic data generation, the prompt template is shown in Figure 6. Figure 7 shows the instruction used during both training and inference.
Figure 6: The prompt we used for synthetic data generation.
Figure 7: Instruction prompt used during both training and inference for our guardrail.
C.2 Human evaluation
We validated our LLM-generated synthetic translations (Section 3) by comparing the model’s safety judgments with human judgments. We have subsampled 100 sentences for ZH, RU, and AR. For each language, we have one volunteer, and the volunteers are native speakers in the target languages and are proficient in English. The guidelines of human evaluation is shown in Figure 8, and the results are listed in Table 9. The safety preservation rate is a three-point scale measurement (0-2). 0: The safety label is not preserved; 1: the safety label is preserved, but meaning is altered; 2: both safety label and original meaning are preserved.
Lang AR ZH RU
Translation 1.98 1.75 1.9
Local 1.67 1.73 1.91
Challenging 1.32 1.58 1.75
Table 9: Here we report human annotation on safety preservation rate of synthetic data generation. Higher score the better. The scale is from 0 to 2.
Figure 8: Guidelines for human annotation of synthetic multilingual training data.
Appendix D Multilingual Variants
Following (Yoo et al., 2024), we generate code-switching prompts to evaluate the guardrail’s robustness to multilingual variations. Rather than mixing 10 different languages, we construct prompts by combining English with a single additional language, reflecting a more realistic code-switching scenario in practical applications. We generate code-switching prompts for MultiJail and XSafety dataset.
Additionally, following (Upadhayay and Behzadan, 2024), we observe that LLMs are more likely to produce harmful responses when distractors from low-resource languages are present. Specifically, we embed jailbreaking prompts between benign questions written in a low-resource language. In our experiments, jailbreaking prompts are drawn from the MultiJail and XSafety datasets, while benign questions are selected from the Vietnamese subset of the XQuAD dataset (Artetxe et al., 2019). Figure 9 illustrates examples of the generated code-switching (CSRT) and sandwich attacks used to evaluate multilingual guardrail robustness.
Figure 9: Examples of code-switching (CSRT) and sandwich attack (sandwich) prompts for multilingual robustness evaluation.
Appendix E GRPO algorithm
The objective function of GRPO is defined as follow. Let $\pi_{\theta}$ denote the language model parameterized by $\theta$, and $\pi_{\text{old}}$ be the model from a previous iteration. Given a prompt $p$, we sample a group of generations $\{o_{1},\ldots,o_{G}\}\sim\pi_{\text{old}}(\cdot\mid p)$, each associated with a reward $\{r_{1},\ldots,r_{G}\}$. Let $\mathbb{D}_{\text{KL}}$ denote the KL-divergence between two distributions. GRPO estimates the advantage of a generation $o_{i}$ using:
$A_{i}=\frac{r_{i}-\operatorname{mean}(\{r_{1},\ldots,r_{G}\})}{\operatorname{std}(\{r_{1},\ldots,r_{G}\})}.$
The GRPO objective is then defined as:
$\mathcal{J}_{\rm GRPO}$ $=\frac{1}{G}\sum_{i=1}^{G}[\min(\frac{\pi_{\theta}(o_{i}|p)}{\pi_{\theta{\rm old}}{(o_{i}|p)}}A_{i},$
$clip(\frac{\pi_{\theta}(o_{i}|p)}{\pi_{\theta{\rm old}}{(o_{i}|p)}},1-\epsilon,1+\epsilon)A_{i}))]$
$-\beta\mathbb{D}_{\rm KL}[\pi_{\theta}||\pi_{\rm ref}].$
Appendix F Additional Results on QWEN
Table 10 shows additional results from training QWEN2.5-3B (Qwen et al., 2025) with our framework.
F1 RTP_LX Aya XSafety MultiJail
QWEN 89.76 98.26 95.43 91.56
Table 10: Here we report F1 scores of QWEN-2.5-3B across different datasets. We take the average across both in-domain and out-of-domain languages.
Appendix G Additional Results
We here show detailed break-down results on additional datasets.
Figure 10: F1 score breakdown on the XSafety dataset, evaluated across 8 target languages.
Figure 11: F1 score breakdown on the aya dataset, evaluated across 8 target languages.
Figure 12: F1 score breakdown on the PTP_wildchat dataset, evaluated across 8 target languages.
Figure 13: F1 score breakdown on the Multijail dataset, evaluated across 8 target languages.
We additionally perform code-switching and sandwich attack on XSafety dataset and the results are shown in Table 11 and Table 12.
Models EN $\uparrow$ Avg-CSRT $\uparrow$ $\Delta$downarrow
DUO-Guard 75.71 70.19 5.52
GuardR 85.99 84.42 1.57
LlaMa-Guard-3 64.48 61.08 3.40
Aegis-2.0 66.22 39.30 26.92
Wildguard 91.40 84.40 7.00
MrGuard 92.71 91.51 1.21
Table 11: F1 scores on code-switching prompts evaluated on the XSafety datasets. The best-performing results across models are highlighted in bold. $\Delta$ represents the difference between the F1 score on English prompts and the averaged F1 score over all code-switching variants across both ID and OOD languages.
Models
Avg-
Orig
$\uparrow$
Avg-
Sandwich
$\uparrow$
$\Delta$$\downarrow$
DUO-Guard 64.77 1.23 63.53
GuardR 82.25 74.83 7.42
LlaMa-Guard-3 61.50 6.37 55.13
Aegis-2.0 35.32 1.12 34.20
Wildguard 69.21 42.38 26.83
MrGuard 93.48 81.13 12.36
Table 12: F1 scores on sandwich attacks evaluated on the XSafety datasets. The best-performing results across models are highlighted in bold. Avg-Orig indicates the average F1 score on before attack, and the average F1 score after sandwich attack across both ID and OOD languages. $\Delta$ represents the difference between them.
Moreover, instead of using a curriculum-based language reward, we can assign a fixed reward to promote multilingual reasoning. In our experiments, we set a constant language reward of 0.5 for all non-English explanations (see Table 13).
RTP_LX Aya Xsafety Wildchat Multijail
Curr 89.27 98.18 93.48 91.54 96.50
Fixed 89.13 98.22 93.02 91.46 96.30
Table 13: Comparison between fixed and curriculum-based language rewards. Here we report average F1 score across various languages on five datasets.Curriculum refers to the reward as described in Section 3.3 while Fixed denotes a constant reward applied to all non-English reasoning.
G.1 Qualitative Result
We provided some qualitative results generated from our multilingual guardrail in Figure 14.
Figure 14: Example of reasoning generated from our multilingual guardrail.